# Probing quantum complexity via universal saturation of stabilizer entropies

### v1

@@ I Introduction
Nonstabilizerness or ‘magic’ has been proposed as the resource theory of fault-tolerant quantum computers [1]. It lower bounds the non-Clifford resources needed to run quantum computers [2] and relates to the complexity of classical simulation algorithms based on Clifford operations [3].
@@
Nonstabilizerness monotones are non-increasing under Clifford operations, while applying non-Clifford operations can enhance the amount of nonstabilizerness of a state. On a quantitative level, the relationship between nonstabilizerness and number of non-Clifford operations is not well understood [4, 5]. As nonstabilizerness is a measure of complexity for quantum computers and a necessary condition for quantum advantage, it is paramount to understand how nonstabilizerness depends on the number of non-Clifford operations [6, 7, 8, 4].
@@
A wide range of magic monotones have been proposed, such as stabilizer rank [3], min-relative entropy of magic [9, 4] and log-free robustness of magic [2, 4]. These measures probe different aspects of nonstabilizerness. In particular, the log-free robustness of magic can be related to the complexity of a classical simulation algorithm, while the min-relative entropy of magic is related to the distance to the closest stabilizer state. However, they require an optimization program to be computed, making them in general intractable for the study of larger system sizes. Thus, methods to feasibly approximate these monotones are desired.
@@
$\alpha$-Stabilizer Rényi entropies (SREs) have been proposed to quantitatively explore nonstabilizerness harnessing their efficient numerical [10, 11, 12], analytic [13] and experimental accessibility [14, 15, 16, 17]. Here, the Rényi index $\alpha$ indicates the moment of the SRE. They are monotones for pure states for $\alpha\geq 2$ [18], while monotonicity can be violated for $\alpha<2$ [19]. SREs relate to phases of error-corrected circuits [20], quantify the entanglement spectrum [21], bound fidelity estimation [22, 23, 24], characterize the robustness of shadow tomography [25], and characterize pseudorandom states [26, 27]. Further, SREs characterize many-body phenomena such as phase transitions [11], frustration [28], random matrix product states [29, 30, 31, 32], localization [33] and out-of-time-order correlators [5, 34, 23]. SREs also lower bound other intractable magic monotones [19, 16].
@@
II Stabilizer Rényi entropy
For an $N$-qubit state |$\ket{\psi}$, the $\alpha$-SRE is given by [10]
@@ II Stabilizer Rényi entropy
where $\alpha$ is the index of the SRE and $\mathcal{P}=\{\sigma\}_{\sigma}$ is the set of all $4^{N}$ Pauli strings $\sigma\in\{I,X,Y,Z\}^{N}$ which are tensor products of Pauli matrices. We note the limit $\alpha\rightarrow 1$, which can be easily shown using l’Hôpital’s rule and is called the von Neumann stabilizer entropy [19]
@@
$M_{\alpha}$ is a resource measure of nonstabilizerness for pure states [1]. SREs are faithful, i.e. $M_{\alpha}=0$ only for pure stabilizer states, while $M_{\alpha}>0$ for all other states. Further, $M_{\alpha}$ is a monotone for $\alpha\geq 2$ [18], i.e. non-increasing under free operations, which are Clifford operations that map pure states to pure states. Note that for $\alpha<2$ SREs can violate the monotonicity condition [19]. For all $\alpha$, SREs are invariant under Clifford unitaries $U_{\text{C}}$, i.e. $M_{\alpha}(U_{\text{C}}\ket{\psi})=M_{\alpha}(\ket{\psi})$. SREs are also additive with $M_{\alpha}(\ket{\psi}\otimes\ket{\phi})=M_{\alpha}(\ket{\psi})+M_{\alpha}(\ket{\phi})$. Note that $M_{\alpha}$ are not strong monotones for any $\alpha$ [19]. As Rényi entropies, SREs are monotonously increasing with decreasing $\alpha$, i.e.
@@
As other nonstabilizerness monotones, we also consider the min-relative entropy of magic [4]
@@
where the maximum is taken over the set of pure stabilizer states. $D_{\text{min}}$ measures the distance between |$\ket{\psi}$ and its closest stabilizer state. It is upper bounded by $D_{\text{min}}\leq N\ln(2)$, which is asymptotically reached for Haar random states [4]. The SRE lower bounds $D_{\text{min}}$ for $N>1$
@@
We also consider the log-free robustness of magic [2, 4]
@@
where $S=\{\eta_{i}\}_{i}$ is the set of pure $N$-qubit stabilizer states. The SRE lower bounds the log-free robustness for $\alpha\geq 1/2$ [2, 10]
@@
III Clifford circuits with T-gates
We now study the SRE for random circuits composed of Clifford unitaries and T-gates. We consider a circuit of $N_{\text{T}}$ layers consisting of randomly sampled Clifford circuits $U_{\text{C}}^{(k)}$ and the single-qubit T-gate $T=\text{diag}(1,\exp(-i\pi/4)$
@@ III Clifford circuits with T-gates
For $N_{\text{T}}=0$, we have Clifford states, while for $N_{\text{T}}\sim N$ we have highly random states [5].
@@
Analytic SRE. The average SRE of such states is known exactly for $\alpha=2$ [10]
@@
where we defined the $T$-gate density $q=N_{\text{T}}/N$. We study the SRE density $m_{2}=M_{2}/N$ in Fig.2a for different $N$. We observe that $m_{2}$ increases linearly with $q$ and converges to a constant for large $q$. For large $N$, we observe that the convergence appears to be a sharp transition to the maximal SRE $m_{2}=\ln(2)$ which occurs at a critical T-gate density $q_{\text{c},2}$. We determine $q_{\text{c},2}$ by studying the scaling at finite $N$ [35]. We find that $\partial_{q}m_{2}$ as function of $q$ exhibits scale-invariant properties, i.e. the curves for different $N$ can be mapped onto each other by appropriate rescaling around the critical point, a hallmark for phase transitions [35]. In particular, we find that $\partial_{q}m_{2}$ intersects for all $N$ at the same point, which gives us $q_{\text{c},2}$. Using Eq. (12), we find that the intersection occurs for all $N$ at the critical T-gate density
@@
In Fig.2b we plot the derivative $\partial_{q}m_{2}$, observing that for all $N$ the curves indeed intersect at $q_{\text{c},2}$. In Fig.2c, we observe that by shifting $q$ with $q_{\text{c},2}$ and rescaling with $N$, we can collapse the curves of different $N$, as expected for the scale-invariant behavior close to critical points [35].
@@
$M_{0}(\ket{\psi})=\ln(\sum_{\sigma}\Theta(\bra{\psi}\sigma\ket{\psi}^{2}))-N\ln(2)$ (14)
where $\Theta(x)$ is the Heaviside function with $\Theta(x)=0$ for $x\leq 0$ and $\Theta(x)=1$ for $x>0$. Stabilizer states are stabilized by a commuting subgroup $G$ of $2^{N}$ Pauli strings with |$\ket{\psi}\sigma\ket{\psi}^{2}=1$ for $\sigma\in G$. The group $G$ has $N$ generators. Applying a T-gate on a stabilizer state breaks at most one generator of $G$, resulting in a state with $N-1$ generators and $2^{N-1}$ stabilizing Pauli strings. In fact, each additional T-gate can break only one generator. Thus, we find $M_{0}\leq N_{\text{T}}\ln(2)$ and $M_{0}\leq N\ln(2)$ [10]. When the T-gate is applied after a random Clifford circuit, the $T$-gate will break one of the generators of $G$ with overwhelming probability. Thus, with overwhelming probability $N_{\text{T}}\approx N$ T-gates are necessary and sufficient to reach $M_{0}=0$, thus the critical T-gate density is $q_{\text{c},0}=1$.
@@
Next, we estimate the value of the maximal SRE for $N\gg 1$. First, we recall the fact that nonstabilizerness approaches the asymptotically the maximal possible value for randomly chosen states [4]. Thus, by estimating the SRE of a random state for large $N$ we can approximate $M_{\alpha}^{\text{max}}$. A random state is spread out over the whole Pauli spectrum where for simplicity we assume a uniform distribution with $\bra{\psi}\sigma\ket{\psi}^{2}=2^{-N}$ for $\sigma\in\mathcal{P}/\{I\}$. While such a state is not a positive density matrix, we find from numerical studies that this spectrum is a sufficiently good approximation of an actual random Pauli spectrum for large $N$. Using this ansatz, we find
@@
Now, in the limit $N\gg 1$ we find
@@
IV Random basis evolution
We proceed to investigate another circuit model which can be seen as a type of random time evolution. It consist of a deep Clifford circuit as in Eq. (10) with many layers $d$, but replace the T-gates with a parameterized rotation $R_{z}(\theta)=\exp(-i\theta/2\sigma_{z})$. A similar model with randomly chosen $\theta$ has been shown to produce highly random states [36]. Here, instead we choose very small $\theta$. In particular, we choose $\theta=2t/\sqrt{d}$ with $d\gg N$
@@ IV Random basis evolution
For $t=0$, Eq. 19 gives a Clifford state |$\ket{0}$, while for $t\sim\sqrt{N}$ the SRE converges to the average value of the SRE for Haar random states. We compute $M_{2}$ for Eq. (19) analytically using the result of Ref. [10]
@@
up to a time $t_{\text{c}}^{2}\approx N\ln(2)$, where it then converges to the average value of Haar random states (see Appendix B). From the bound Eq. (6) we have for $\alpha=2$ [19]
@@
V Random Hamiltonian evolution
Next, we study the evolution of states under random Hamiltonians [37]. We evolve an initial random stabilizer state |$\ket{\psi(0)}=\ket{\psi_{\text{STAB}}}$ state
@@ V Random Hamiltonian evolution
Due to Levy’s lemma, observed expectation values such as $F(t)$ concentrate with exponentially high probability around its average for each sampled state [38].
@@
Now, how does the Pauli spectrum $\beta_{\sigma}(t)^{2}$ change when the stabilizer state is evolved in time $t$? For $t=0$, the Pauli spectrum has two peaks at $\beta_{\sigma}^{2}=0$ and $\beta_{\sigma}^{2}=1$. For $t>0$, the two peaks shift and diffuse. However, we observe numerically that the two peaks remain highly concentrated even for relatively large $t$. This can be observed in Fig. 5a, where we plot the histogram of the Pauli spectrum. Note that up to $t\lesssim 1$, there are two distinct peaks with a gap in between them. Note that Fig. 5a is a logarithmic plot, and the peak for small $\beta_{\sigma}^{2}$ appears broad in logarithmic space, but is actually very concentrated close to its mean value. Let us now approximate the two peaks as Delta-functions centered around their mean value. For many qubits $N\gg 1$, we can easily compute the average of each peak, i.e. $\beta_{\sigma\in G}^{2}\approx F^{2}$, and $\beta_{\sigma\notin G}^{2}\approx 2^{-N}(1-F^{2})$.

### v2

@@ 1 Introduction
Nonstabilizerness or ‘magic’ has been proposed as the resource theory of fault-tolerant quantum computers [Veitch et al.(2014)Veitch, Mousavian, Gottesman, and Emerson]. It lower bounds the non-Clifford resources needed to run quantum computers [Howard and Campbell(2017)] and relates to the complexity of classical simulation algorithms based on Clifford operations [Bravyi et al.(2016)Bravyi, Smith, and Smolin].
@@
Nonstabilizerness monotones are non-increasing under Clifford operations, while applying non-Clifford operations can enhance the amount of nonstabilizerness of a state. On a quantitative level, the relationship between quantum complexity and number of non-Clifford operations is an intensely studied topic [Haferkamp et al.(2022)Haferkamp, Montealegre-Mora, Heinrich, Eisert, Gross, and Roth, Liu and Winter(2022), Leone et al.(2021)Leone, Oliviero, Zhou, and Hamma]. As nonstabilizerness is a measure of complexity for quantum computers and a necessary condition for quantum advantage, it is paramount to understand how nonstabilizerness depends on the number of non-Clifford operations [White et al.(2021)White, Cao, and Swingle, Ellison et al.(2021)Ellison, Kato, Liu, and Hsieh, Sarkar et al.(2020)Sarkar, Mukhopadhyay, and Bayat, Liu and Winter(2022)].
@@
A wide range of magic monotones have been proposed, such as stabilizer rank [Bravyi et al.(2016)Bravyi, Smith, and Smolin], min-relative entropy of magic [Bravyi et al.(2019)Bravyi, Browne, Calpin, Campbell, Gosset, and Howard, Liu and Winter(2022)] and log-free robustness of magic [Howard and Campbell(2017), Liu and Winter(2022)]. These measures probe different aspects of nonstabilizerness. In particular, the log-free robustness of magic can be related to the complexity of a classical simulation algorithm, while the min-relative entropy of magic is related to the distance to the closest stabilizer state. However, they require an optimization program to be computed, making them in general intractable for the study of larger system sizes. Thus, methods to feasibly approximate these monotones are desired.
@@
$\alpha$-Stabilizer Rényi entropies (SREs) have been proposed to quantitatively explore nonstabilizerness harnessing their efficient numerical [Leone et al.(2022)Leone, Oliviero, and Hamma, Haug and Piroli(2023a), Lami and Collura(2023)], analytic [López and Kos(2024), Turkeshi et al.(2023a)Turkeshi, Dymarsky, and Sierant] and experimental accessibility [Haug and Kim(2023), Oliviero et al.(2022)Oliviero, Leone, Hamma, and Lloyd, Haug et al.(2024)Haug, Lee, and Kim, Bluvstein et al.(2022)Bluvstein, Levine, Semeghini, Wang, Ebadi, Kalinowski, Keesling, Maskara, Pichler, Greiner, et al.]. Here, the Rényi index $\alpha$ indicates the moment of the SRE. They are monotones for pure states for $\alpha\geq 2$ [Leone and Bittel(2024)], while monotonicity can be violated for $\alpha<2$ [Haug and Piroli(2023b)]. SREs relate to phases of error-corrected circuits [Niroula et al.(2023)Niroula, White, Wang, Johri, Zhu, Monroe, Noel, and Gullans], quantify the entanglement spectrum [Tirrito et al.(2024)Tirrito, Tarabunga, Lami, Chanda, Leone, Oliviero, Dalmonte, Collura, and Hamma], bound fidelity estimation [Leone et al.(2024a)Leone, Oliviero, Esposito, and Hamma, Leone et al.(2023)Leone, Oliviero, and Hamma, Hinsche et al.(2024)Hinsche, Ioannou, Jerbi, Leone, Eisert, and Carrasco], characterize the robustness of shadow tomography [Brieger et al.(2023)Brieger, Heinrich, Roth, and Kliesch] and characterize pseudorandom states [Gu et al.(2024a)Gu, Leone, Ghosh, Eisert, Yelin, and Quek, Haug et al.(2023)Haug, Bharti, and Koh]. Further, SREs characterize many-body phenomena [Gu et al.(2024b)Gu, Oliviero, and Leone] such as phase transitions [Haug and Piroli(2023a), Tarabunga et al.(2023)Tarabunga, Tirrito, Chanda, and Dalmonte], frustration [Odavić et al.(2023)Odavić, Haug, Torre, Hamma, Franchini, and Giampaolo], random matrix product states [Frau et al.(2024)Frau, Tarabunga, Collura, Dalmonte, and Tirrito, Lami et al.(2024)Lami, Haug, and De Nardis, Paviglianiti et al.(2024)Paviglianiti, Lami, Collura, and Silva, Mello et al.(2024)Mello, Santini, and Collura], localization [Turkeshi et al.(2023b)Turkeshi, Schirò, and Sierant] and out-of-time-order correlators [Leone et al.(2021)Leone, Oliviero, Zhou, and Hamma, Garcia et al.(2023)Garcia, Bu, and Jaffe, Leone et al.(2023)Leone, Oliviero, and Hamma]. SREs also lower bound other intractable magic monotones [Haug and Piroli(2023b), Haug et al.(2024)Haug, Lee, and Kim].
@@
2 Stabilizer Rényi entropy
For an $N$-qubit state |$\ket{\psi}$, the $\alpha$-SRE is given by [Leone et al.(2022)Leone, Oliviero, and Hamma]
@@ 2 Stabilizer Rényi entropy
where $\alpha$ is the index of the SRE and $\mathcal{P}=\{\sigma\}_{\sigma}$ is the set of all $4^{N}$ Pauli strings $\sigma\in\{I,X,Y,Z\}^{N}$ which are tensor products of Pauli matrices. We note the limit $\alpha\rightarrow 1$, which can be easily shown using l’Hôpital’s rule and is called the von Neumann stabilizer entropy [Haug and Piroli(2023b)]
@@
$M_{\alpha}$ is a resource measure of nonstabilizerness for pure states [Veitch et al.(2014)Veitch, Mousavian, Gottesman, and Emerson]. SREs are faithful, i.e. $M_{\alpha}=0$ only for pure stabilizer states, while $M_{\alpha}>0$ for all other states. Further, $M_{\alpha}$ is a monotone for $\alpha\geq 2$ [Leone and Bittel(2024)], i.e. non-increasing under free operations, which are Clifford operations that map pure states to pure states. Note that for $\alpha<2$ SREs can violate the monotonicity condition [Haug and Piroli(2023b)]. For all $\alpha$, SREs are invariant under Clifford unitaries $U_{\text{C}}$, i.e. $M_{\alpha}(U_{\text{C}}\ket{\psi})=M_{\alpha}(\ket{\psi})$. SREs are also additive with $M_{\alpha}(\ket{\psi}\otimes\ket{\phi})=M_{\alpha}(\ket{\psi})+M_{\alpha}(\ket{\phi})$. Note that $M_{\alpha}$ are not strong monotones for any $\alpha$ [Haug and Piroli(2023b)]. As Rényi entropies, SREs are monotonously increasing with decreasing $\alpha$, i.e.
@@
As other nonstabilizerness monotones, we also consider the min-relative entropy of magic [Bravyi et al.(2019)Bravyi, Browne, Calpin, Campbell, Gosset, and Howard, Liu and Winter(2022)]
@@
where the maximum is taken over the set of pure stabilizer states. Here, $F_{\text{STAB}}=\max_{\ket{\phi}\in\text{STAB}}|\braket{\psi}{\phi}|^{2}$ is the stabilizer fidelity. $D_{\text{min}}$ measures the distance between |$\ket{\psi}$ and its closest stabilizer state. It is upper bounded by $D_{\text{min}}\leq N\ln(2)$, which is asymptotically reached for Haar random states [Liu and Winter(2022)]. The SRE lower bounds $D_{\text{min}}$ for $N>1$
@@
We also consider the log-free robustness of magic [Howard and Campbell(2017), Liu and Winter(2022)]
@@
where $S=\{\eta_{i}\}_{i}$ is the set of pure $N$-qubit stabilizer states. The SRE lower bounds the log-free robustness for $\alpha\geq 1/2$ [Howard and Campbell(2017), Leone et al.(2022)Leone, Oliviero, and Hamma]
@@
3 Clifford circuits with T-gates
We now study the SRE for random circuits composed of Clifford unitaries and T-gates [Haferkamp et al.(2022)Haferkamp, Montealegre-Mora, Heinrich, Eisert, Gross, and Roth]. We consider a circuit of $N_{\text{T}}$ layers consisting of randomly sampled Clifford circuits $U_{\text{C}}^{(k)}$ and the single-qubit T-gate $T=\text{diag}(1,\exp(-i\pi/4)$
@@ 3 Clifford circuits with T-gates
For $N_{\text{T}}=0$, we have Clifford states, while for $N_{\text{T}}\sim N$ we have highly random states [Leone et al.(2021)Leone, Oliviero, Zhou, and Hamma].
@@
Analytic SRE. The average SRE of such states is known exactly for $\alpha=2$ [Leone et al.(2022)Leone, Oliviero, and Hamma]
@@
where we defined the $T$-gate density $q=N_{\text{T}}/N$. We study the SRE density $m_{2}=M_{2}/N$ in Fig.2a for different $N$. We observe that $m_{2}$ increases linearly with $q$ and converges to a constant for large $q$. For large $N$, we observe that the convergence appears to be a sharp transition to the maximal SRE $m_{2}=\ln(2)$ which occurs at a critical T-gate density $q_{\text{c},2}$. We determine $q_{\text{c},2}$ by studying the scaling at finite $N$ [Osterloh et al.(2002)Osterloh, Amico, Falci, and Fazio]. We find that $\partial_{q}m_{2}$ as function of $q$ exhibits scale-invariant properties, i.e. the curves for different $N$ can be mapped onto each other by appropriate rescaling around the critical point, a hallmark for phase transitions [Osterloh et al.(2002)Osterloh, Amico, Falci, and Fazio]. In particular, we find that $\partial_{q}m_{2}$ intersects for all $N$ at the same point, which gives us $q_{\text{c},2}$. Using Eq. (12), we find that the intersection occurs for all $N$ at the critical T-gate density
@@
In Fig.2b we plot the derivative $\partial_{q}m_{2}$, observing that for all $N$ the curves indeed intersect at $q_{\text{c},2}$. In Fig.2c, we observe that by shifting $q$ with $q_{\text{c},2}$ and rescaling with $N$, we can collapse the curves of different $N$, as expected for the scale-invariant behavior close to critical points [Osterloh et al.(2002)Osterloh, Amico, Falci, and Fazio].
@@
$M_{0}(\ket{\psi})=\ln\Big{(}\sum_{\sigma}\Theta(\bra{\psi}\sigma\ket{\psi}^{2})\Big{)}-N\ln(2)$ (14)
where $\Theta(x)$ is the Heaviside function with $\Theta(x)=0$ for $x\leq 0$ and $\Theta(x)=1$ for $x>0$. Stabilizer states are stabilized by a commuting subgroup $G$ of $2^{N}$ Pauli strings with |$\ket{\psi}\sigma\ket{\psi}^{2}=1$ for $\sigma\in G$. The group $G$ has $N$ generators. Applying a T-gate on a stabilizer state breaks at most one generator of $G$, resulting in a state with $N-1$ generators and $2^{N-1}$ stabilizing Pauli strings. In fact, each additional T-gate can break only one generator. Thus, we find $M_{0}\leq N_{\text{T}}\ln(2)$ and $M_{0}\leq N\ln(2)$ [Leone et al.(2022)Leone, Oliviero, and Hamma]. When the T-gate is applied after a random Clifford circuit, the $T$-gate will break one of the generators of $G$ with overwhelming probability. Thus, with overwhelming probability $N_{\text{T}}\approx N$ T-gates are necessary and sufficient to reach $M_{0}=0$, thus the critical T-gate density is $q_{\text{c},0}=1$.
@@
Next, we estimate the value of the maximal SRE for $N\gg 1$. First, we recall the fact that nonstabilizerness approaches the asymptotically the maximal possible value for randomly chosen states [Liu and Winter(2022)]. Thus, by estimating the SRE of a random state for large $N$ we can approximate $M_{\alpha}^{\text{max}}$. A random state is spread out over the whole Pauli spectrum where for simplicity we assume a uniform distribution with $\bra{\psi}\sigma\ket{\psi}^{2}=2^{-N}$ for $\sigma\in\mathcal{P}/\{I\}$. While such a state is not a positive density matrix, we find from numerical studies that this spectrum is a sufficiently good approximation of an actual random Pauli spectrum for large $N$. Using this ansatz, we find
@@
This result is confirmed by the average $M_{\alpha}(\ket{\psi_{\text{Haar}}})$ for Haar random states which has been computed in Ref. [Turkeshi et al.(2023a)Turkeshi, Dymarsky, and Sierant], where Haar random states are known to have asymptotically maximal nonstabilizerness [Liu and Winter(2022)]. Now, in the limit $N\gg 1$ we find
@@
4 Random basis evolution
We proceed to investigate another circuit model which can be seen as a type of random time evolution. It consist of a deep Clifford circuit as in Eq. (10) with many layers $d$, but replace the T-gates with a parameterized rotation $R_{z}(\theta)=\exp(-i\theta/2\sigma_{z})$. A similar model with randomly chosen $\theta$ has been shown to produce highly random states [Haah et al.(2024)Haah, Liu, and Tan]. Here, instead we choose very small $\theta$. In particular, we choose $\theta=2t/\sqrt{d}$ with $d\gg N$
@@ 4 Random basis evolution
For $t=0$, Eq. 19 gives a Clifford state |$\ket{0}$, while for $t\sim\sqrt{N}$ the SRE converges to the average value of the SRE for Haar random states. We compute $M_{2}$ for Eq. (19) analytically using the result of Ref. [Leone et al.(2022)Leone, Oliviero, and Hamma]
@@
up to a time $t_{\text{c}}^{2}\approx N\ln(2)$, where it then converges to the average value of Haar random states (see Appendix B). From the bound Eq. (6) we have for $\alpha=2$ [Haug and Piroli(2023b)]
@@
5 Random Hamiltonian evolution
Next, we study the evolution of states under random Hamiltonians [Cotler et al.(2017)Cotler, Hunter-Jones, Liu, and Yoshida]. We evolve an initial random stabilizer state |$\ket{\psi(0)}=\ket{\psi_{\text{STAB}}}$ state
@@ 5 Random Hamiltonian evolution
Due to Levy’s lemma, observed expectation values such as $F(t)$ concentrate with exponentially high probability around its average for each sampled state [Popescu et al.(2006)Popescu, Short, and Winter].
@@
Now, how does the Pauli spectrum $\beta_{\sigma}(t)^{2}$ change when the stabilizer state is evolved in time $t$? For $t=0$, the Pauli spectrum has two peaks at $\beta_{\sigma}^{2}=0$ and $\beta_{\sigma}^{2}=1$. For $t>0$, the two peaks shift and diffuse. However, we observe numerically that the two peaks remain highly concentrated even for relatively large $t$. Note that for Haar random states, the concentration has been proven [Turkeshi et al.(2023a)Turkeshi, Dymarsky, and Sierant]. We show the histogram of the Pauli spectrum in Fig. 5a Note that up to $t\lesssim 1$, there are two distinct peaks with a gap in between them. Note that Fig. 5a is a logarithmic plot, and the peak for small $\beta_{\sigma}^{2}$ appears broad in logarithmic space, but is actually very concentrated close to its mean value. Let us now approximate the two peaks as Delta-functions centered around their mean value. For many qubits $N\gg 1$, we can easily compute the average of each peak, i.e. $\beta_{\sigma\in G}^{2}\approx F^{2}$, and $\beta_{\sigma\notin G}^{2}\approx 2^{-N}(1-F^{2})$.